# Branching Flows: Discrete, Continuous, and Manifold Flow Matching with Splits and Deletions

### v1

@@ 3.8 Concrete Schemes for Sampling $Z$
Here we outline concrete schemes for sampling $Z:=(X_{1}^{+ø},X_{0},\mathcal{T},\mathcal{A})$, starting from $X_{1}$, that we use in our empirical examples below. Some datasets have elements that are naturally grouped (e.g. proteins, grouped by chain), and one can construct $Z$ to respect these groupings, where it makes sense to do so. In the case where there is only one group, all elements have the same group index. Further, if any elements are ‘fixed’ (e.g. in the protein example below, by a conditioning mask), then any elements with a fixed element between them in the sequence must belong to a different group.
3.8.1 Initial Distribution: $X_{0}\sim\mathcal{D}_{0}|(X_{1}=x_{1})$
@@
3.8.3 Forest: $\mathcal{T}\sim\mathcal{D}_{\mathcal{T}}|(X_{1}^{+ø}=x_{1}^{+ø},X_{0}=x_{0})$
@@
3.8.4 Internal Anchors: $\mathcal{A^{I}}\sim\mathcal{D}|(X_{1}^{+ø}=x_{1}^{+ø},X_{0}=x_{0},\mathcal{T})$
From the root to the leaves of the tree, the elements are drifting towards a sequence of anchors, which change each time a split event occurs. Thus care must be taken when sampling the anchors to maximize the learnability of the generator. One useful approach is to specify an anchor merging distribution $p(a_{\text{parent}}|a_{\text{child}_{1}},a_{\text{child}_{2}},\mathcal{T})$ and recursively sample the anchors from leaves to root. Since ordered sequences of continuous values (especially for spatial positions) often exhibit strong autocorrelation with similar values for neighbouring elements, but discrete values typically do not, the anchor merging schemes differ depending on whether the state components are continuous (including manifolds) or discrete:
  • Continuous: Geodesic interpolation between $a_{\text{child}_{1}}$ and $a_{\text{child}_{2}}$, optionally weighted by the number of descendants of each child.
@@
4 Results
We consider three datasets: i) small molecules, where we jointly generate atom positions and atom labels, ii) Antibody heavy chains (discrete sequence only), and iii) protein backbones, comprising amino acid ‘frames’ (jumper2021highly) with Euclidean positions & SO(3) orientations, and discrete amino acid sequence labels.
We are interested in understanding the degree to which Branching Flows-trained models can learn the data distribution. A strategy we use for QM9 and antibody heavy chains is to consider a set of domain-specific projections from the sample space to a scalar dimension, and then compare 10,000 Branching Flows samples to 10,000 samples from the data distribution. We also use UMAP (mcinnes2020umapuniformmanifoldapproximation) (or seqUMAP, hanke2022multivariate for the antibody sequence case) for projections into two dimensions which can then be used to visualize the sample and data distribution.
4.1 QM9
Figure 3: QM9 Branching Flows sampling trajectories. A visualization of inference trajectories from a QM9-trained Branching Flows model. The final sampled molecule, when $t=1$, is depicted on the right. At every inference step from $t=0$ to $t=1$, we draw the current $x_{t}$ state as colored points, with a small rightward displacement, which shows the branching and deletion history as $x_{t}$ is transported from a single element when $t=0$ to the final sampled molecule. The color of the trails shows the atom type which begins as ‘Mask’ (in orange) and switches to concrete atoms as $t\rightarrow 1$. For the purposes of visualization we use 2000 steps with a modified step schedule: $t=(1-(\cos(s\pi/2)))$, for $s\in(0,1/2000,2/2000,...,1).$
Figure 4: QM9 data vs generated samples. 10,000 random samples from the QM9 dataset vs 10,000 samples generated using a Branching Flows model. Panel A computes molecular fingerprints (using RDKit) and jointly embeds QM9 data and generated samples using UMAP. Left shows QM9-only embeddings, colored by the number of carbon atoms in each molecule; right shows generated samples, similarly colored, and center shows the overlap of QM9 (blue) vs Branching Flows generated samples (red), showing similar distributions. Panel B depicts univariate property distributions from QM9 and Branching Flows generated samples, including atom counts, as well as a number of standard molecular descriptors computed with RDKit.
QM9 (ramakrishnan2014quantum) contains small organic molecules selected from the GDB (Generated Database) chemical space enumeration. For each molecule, the 3D geometry was relaxed with density functional theory, and a consistent set of quantum mechanical properties were computed. Here we use Branching Flows to build a generative model over the atom coordinates and discrete labels (i.e. we do not consider the molecule properties).
4.1.1 Data
4.1.2 Branching Flows specification
  • Conditional base process:
    • Continuous: OU process flow (Appendix C), with a mean-reversion rate $\theta=5$, and where the variance decays from 10 when $t=0$ to $0.001$ by $t=1$.
    • Discrete: DFM convex interpolation of $x_{0}$, uniform noise, and $x_{1}$, with hazards $\mathcal{D}_{1}=\text{Beta}(2,2),\mathcal{D}_{2}=\text{Beta}(2,2),\omega_{u}=0.2$.
  • Branching hazard distribution: $\operatorname{Beta}(1,3/2)$
  • Coalescence policy: uniformly merge adjacent nodes.
  • Deletion hazard distribution: $U(0,1)$
  • $x_{0}$ length distribution: 1, always.
4.1.3 Model and Training
We use a 12 layer transformer model (12 heads, each with head dimension of 64), with an embedding dimension of 384. The atom positions are encoded by Random Fourier Features (RFFs), and pairwise spatial features are also provided to each layer by a learnable, layer-specific, head-specific attention bias. Rotary Positional Encoding (RoPE, su2023roformerenhancedtransformerrotary) is used to encode the primary sequence ordering. The final six layers additively update the atom positions (and the spatial pair features are recalculated after each position update). The model outputs the endpoint-predicted (i.e. by $t=1$) locations, logits over the endpoint-predicted discrete state, the log of the per-element expected number of future bifurcations by $t=1$, and a deletion binary logit. The model was trained, using Muon (jordan2024muon) with post-warmup learning rate of $0.005$, for 500k batches (where each batch was 128 randomly-sampled QM9 molecules) and a linear learning-rate cooldown for the last 50k batches.
4.1.4 Results
Figure 3 shows typical inference trajectories for QM9 molecules. Starting from a single $x_{0}$ element with a random position and a masked label, the process undergoes a sequence of bifurcations and deletions governed by the rates output from the model, with the discrete states switching along the trajectory as well, terminating at a generated molecule.
The only relevant benchmark for our purposes is whether the model learns, via Branching Flows, to sample from the data distribution. To investigate this, we generated 10k samples and compared these to 10k QM9 samples (See appendix E.1 for details).
Figure 4 compares generated and data distributions for molecule fingerprints, which encode molecule properties into a fixed-length bit vector, embedded into two-dimensions by UMAP, showing no apparent differences. Also shown are univariate histograms of atom counts and molecule properties, with some subtle departures, most visible in $LogP$ and molecular weight.
We also compare 10,000 samples from a QM9-trained ‘Transdimensional Jump’ diffusion model (Campbell2023TransDimensionalGM). We use a Kolmogorov-Smirnov based statistic ($1-KS_{D}$; see appendix E.1) to quantify agreement between the samples and the data distribution, as shown in table 1. For all univariate metrics, Branching Flows better matches the data distribution. We explore these differences further with visualizations in appendix figure 10.
Model mw logp hbd hba C H O N F atoms
Table 1: Distribution agreement ($1-\mathrm{KS}_{D}$, higher is better) between univariate properties of samples from each model vs. QM9 data. From left to right: Molecular weight, LogP, hydrogen-bond donors, hydrogen-bond acceptors, number of: carbons, hydrogens, oxygens, nitrogens, fluorines, and the total number of atoms.
4.2 Antibody Sequences
To investigate Branching Flows on a pure discrete sequence learning task, we consider the problem of antibody amino acid sequence generation. Antibodies exhibit substantial length variation, most concentrated in the complementarity determining regions (CDRs) which are the primary determinants of antibody-antigen interaction. They originate from a stochastic recombination of variable, diversity, and joining germline gene segments, with non-templated junctions in between, and then further diversify via somatic hypermutation (murphy2016janeway).
Here we compare a Branching Flows model to an ‘oracle-length’ discrete flow matching model, which is identical to the Branching Flows model except that it i) has no splits or deletions, and ii) is initialized with a ‘true’ length sampled from the data distribution.
4.2.1 Data
4.2.2 Branching Flows Specification
  • Conditional base process:
    • Discrete: DFM convex interpolation (appendix B) of $x_{0}$, uniform noise, and $x_{1}$, with hazards $\mathcal{D}_{1}=\text{Beta}(1.5,1.5),\mathcal{D}_{2}=\text{Beta}(2,2),\omega_{u}=0.2$.
  • Branching hazard distribution: $\operatorname{Beta}(1,2)$
  • Coalescence policy: uniformly merge adjacent nodes.
  • Deletion hazard distribution: $U(0,1)$
  • To-be-deleted policy: deletion rate $d_{r}=1.2$
  • $x_{0}$ state distribution: masked tokens.
  • $x_{0}$ length distribution: uniform between 110 and 140.
4.2.3 Model and Training
The models (Branching Flows and oracle-length) were trained using a batch size of 64 resulting in a total of 234 375234\,375 training steps for the 15 million sequence dataset. Muon was used with a 2000 step linear warm-up followed by a cosine decay starting at $10^{-3}$ and decaying to $10^{-7}$.
4.2.4 Results
Figure 5 shows the distributions of statistics (Appendix E.2) from antibody sequences generated by both the Branching Flows and oracle-length models. It also includes the perplexity-over-time plot and 2 seqUMAP (hanke2022multivariate) representations of the 10 00010\,000 generated sequences. Generally, these plots show that distributional overlap with natural sequences is similar between Branching Flows and the oracle model, despite the oracle having additional information. As both models were only trained on 15 million sequences, we anticipate that further training will improve the distributional match. Finally, we plot the sequence-position-specific amino acid frequency distributions for all 20 amino acids in figure 6, which shows that Branching Flows can learn strong position dependence despite not having any prior information about total length.
4.3 Multimodal Branching Flows for Protein Structure and Sequence Generation
Figure 8: Branching Flows protein samples. Top left inset shows pseudoMSA scTM refolding scores, and selected samples are shown (rainbow colored for single chains, and per-chain for dimers). To the right or below each is shown an overlay of the sample and the refolded structure (grey). Specifically, the three worst scTMs are depicted (labelled a, b, and c on the inset and the main panel, showing that the low scTMs are driven by unpredictability due to flexible linkers, or incorrect dimerization prediction.
  • for Euclidean positions, a time-inhomogeneous OU process with mean reversion rate of 100, and a variance of 150 at $t=0$ that decays to 0 as $t\rightarrow 1$.
  • for discrete states, a DFM convex interpolation (appendix B) with a $\mathcal{D}_{1}=\text{Beta}(3.0,1.5),\mathcal{D}_{2}=\text{Beta}(2,2),\omega_{u}=0.2$.
We finetuned two models: one to generate chains (‘BranchChain‘), and one to sample segments of contiguous residues (‘BranchSegment‘). For BranchChain, the designable residues were always complete chains, where the chains were randomly selected for masking during training, so it can generate chains unconditionally, or it can generate chains conditioned on other fixed chains. BranchSegment’s conditioning mask during training was a random number of random-length contiguous segments, so multiple segments can be simultaneously designed during inference.
For BranchChain, we initialize each chain to have $1+\text{Poisson}(20)$ $x_{0}$ elements, each drawn from the ChainStorm base element distribution, and for BranchSegment each designable segment has a single $x_{0}$ element. For sampling ‘to-be-deleted‘ elements, we use $d_{r}=1.2$ for BranchChain and $dr=1.1$ for BranchSegment (reasoning that the more strongly informative conditioning information for BranchSegment might require less exploratory flow paths). Samples were generated with 400 steps: $t=1-(\cos(\pi\tau)+1)/2$ for $\tau\in(0,1/400,2/400,...,1).$
For illustration purposes, we also finetuned a BranchChain variant with only a single $x_{0}$ element per chain. Figure 7 shows steps along the sampling trajectory for a sample where two chains (blue and green) were fixed, and two chains were sampled. The orange chain, best seen through the lens of the model-predicted end state $x_{1}^{hat}$, highlights a key Branching Flows feature: at first, the designed chain converges upon two separate domains, with insufficient residues in between them for them to be linked. This is then progressively filled in by insertions, until a complete links connects them as a single chain.
To investigate whether a Branching Flows protein model can generate samples with reasonable protein-like geometry, we sampled, without conditioning, 20 monomers and 20 dimers. Discarding samples with chain breaks (one monomer and five dimers), and samples too short for meaningful evaluation (<25<25 residues in the longest chain: five monomers, and 1 dimer), with 14 monomers and 14 dimers remaining. Using the ‘pseudoMSA refolding’ strategy from ChainStorm (OrestenChainStorm), we use a state-of-the-art protein folding model and test the extent to which it can recover the generated protein geometry from sequences only, which shows that our designs have geometry consistent with the folding model. We generated a pseudoMSA of 100 sequences for each of these with ProteinMPNN (dauparas2022robust), and evaluate the ‘self-consistency TM scores’ of a single refold, via Boltz2 (passaro2025boltz) of this pseudoMSA.
Figure 8 shows the scTM scores, most very close to 1. We also show selected examples, for monomers and dimers, of BranchChain generated proteins and their refolds, specifically showing the three cases with low scTMs. For the two poorly refolded monomers, this is because the generated structure had one or more flexible linkers, and the individual domains were well refolded. For the poorly refolded dimer, this was because BranchChain generated two disconnected chains, and Boltz2 placed these adjacent. Further exploratory sampling suggests that BranchChain does sometimes generate disconnected chains.
Overall, BranchChain generates samples with a variety of protein chain lengths, with plausible backbone geometry. Further, as can be seen in many of the samples in figure 8, the propensity of the ChainStorm base model to sample homodimers is preserved. This is interesting given that ChainStorm’s single attention head that governs symmetric oligomer sampling is primarily driven by chain lengths which are fixed in ChainStorm, but now stochastically generated in BranchChain.
For BranchSegment we do not attempt quantitative evaluations, but we demonstrate that it is a promising solution to the ‘unknown length infix sampling’ problem. Taking a protein complex (PDB: 9IQP, shcheblyakov2025ultra) of a nanobody and the SARS-CoV-2 receptor binding domain (RBD), we mask only the ‘Complementarity Determining Region 3’ (CDR3) as designable, and we generate 20 samples from the RBD-bound complex, and 20 samples from the unbound nanobody alone. The sampled CDR3s had a broad length distribution (7 to 29 residues for bound, 7 to 40 unbound). For the bound samples, even though the side chain atom positions from the RBD were not provided to BranchSegment (it only sees backbone frames and amino acid labels) the designed CDR3s had no atomic clashes with the RBD side chain rotamers for 15 out of 20 of the generated CDR3s. Some of the unbound CDR3s appear structurally plausible, but are extremely unlikely to occur naturally, which is not surprising since neither the base model nor BranchSegment have any antibody or nanobody-specific training.
5 Implementation, Availability, and Key Dependencies
Branching Flows is available as a Julia (Bezanson2017JuliaAF) package BranchingFlows.jl (https://github.com/MurrellGroup/BranchingFlows.jl). This composes with base processes implemented in Flowfusion.jl (https://github.com/MurrellGroup/Flowfusion.jl). Demos and specific models are available via https://github.com/MurrellGroup/BranchingDemos.jl.
We rely on Flux.jl (Flux2018) for our deep learning environment, Zygote.jl for automatic differentiation (Zygote2018), CUDA.jl (besard2018juliagpu) for GPU interaction, Makie.jl (DanischKrumbiegel2021) and Plots.jl (PlotsJL) for visualization, and NNop.jl (https://github.com/pxl-th/NNop.jl) for customized flash attention (dao2022flashattentionfastmemoryefficientexact) kernels. For small molecule processing for the QM9 dataset, we use MoleculeFlow.jl (https://github.com/MoleculeHub/MoleculeFlow.jl), which wraps RDKit (https://www.rdkit.org). Additionally, for the protein model we rely on all the packages listed in OrestenChainStorm.
6 Related Work
Edit Flows
Transdimensional jumps
DrugFlow
7 Discussion
We have shown that Branching Flows is a capable ‘distribution-learner’ of multimodal states that vary in length, addressing a previously-unsolved fundamental problem.
For now, we restrict ourselves to demonstrating new capabilities afforded by Branching Flows, without extensive comparisons on downstream tasks and benchmarks. The reason for this is that performance benefits on benchmarks are often more sensitive to practically useful but theoretically uninteresting domain-specific adjustments to the model and training details, which is orthogonal to the purpose of this manuscript.
For the antibody sequence modeling problem, we show a side-by-side comparison between Branching Flows and an oracle-length model that is provided, via the number of starting tokens, the true length of the sequence. During inference, this length is sampled from the data distribution. We show that both the training dynamics of Branching Flows, and the ability to match the data distribution, are virtually indistinguishable from a model that has been provided the true length.
Note that while training an oracle-length model might be a viable strategy for unconditional generation tasks if any data are available at inference time, sampling the length from the data distribution is not possible for conditional generation tasks (e.g. conditioning on some fixed part of the sequence), as one generally cannot conditionally sample length from the data distribution at test time.
For the QM9 small-molecule dataset, we compare samples from our Branching Flows-trained model with those from the Transdimensional Jump Diffusion model of Campbell2023TransDimensionalGM, showing favorable distribution matching for Branching Flows. We suspect this difference is attributable to two things. Firstly, transdimensional jump diffusion spawns elements at a position that is, in its conditional process, coupled to an element in the data distribution. While the conditional distribution of the spawn point is Gaussian conditioned on the corresponding $X_{1}$ element, the marginal distribution during inference is a (possibly infinite) mixture of Gaussians. Practically, the model must be content with a coarse approximation of this distribution, such that, even in the small step limit, there will be a mismatch between the conditional and marginal paths. Branching Flows avoids this by spawning new elements via a splitting mechanism that, conditioned on $X_{t}$ is independent of $X_{1}$.
In our current implementation, Branching Flows only duplicates elements, but we note that other strategies such as ‘duplicate and add noise’ are also viable, so long as the post-split distribution of the new elements is conditionally independent of $Z$ given $X_{t}$.
Secondly, transdimensional jump diffusion handles discrete atom types by converting them to continuous vectors, which does not compare favorably to discrete character diffusion models (sahoo2024simpleeffectivemaskeddiffusion). Branching Flows allows composition of the branching and deletion process with any underlying flow matching base process, naturally accommodating true multimodal state spaces.
However distribution matching differences may also be attributable to model architecture differences or training details. The transdimensional jump model uses a domain-specific equivariant architecture, while Branching Flows uses a generic one, for example. A side-by-side comparison with identical architecture and training might alter the results, in either direction.
For our protein structure models, we do not perform extensive validation. Most standard protein benchmark metrics are problematic. For now we content ourselves in showing that Branching Flows generates samples that have geometry consistent with a state-of-the-art folding model.
In terms of capabilities, as far as we are aware ‘BranchSegment’ is the first generative model that can perform ‘unknown-length infix sampling’ for proteins, where the backbone is constrained either side of the designable region, and the length of the designed segment is determined by the model. In practice, current generative models can use a range of pre-sampled lengths, with post-sampling filters to select from these, and non-generative models can perform sequence search on a surrogate objective (e.g. via a folding model). Neither will provide distribution-matching guarantees, and both are computationally expensive. Further, the ‘pre-sampled length’ strategy required by fixed-length flow and diffusion models can combinatorically explode if there are multiple sequentially non-contiguous but spatially-interacting segments that need to be designed, which would be the case during protein-protein interface design or re-design.
One practically useful feature, demonstrated with ‘BranchChain’ and ‘BranchSegment’ is that Branching Flows can be finetuned into an existing model. A few days on a single GPU was sufficient. This suggests that, even though the base model was trained via fixed-length flow matching, the internal representations required to learn the branching and deletion process may already be present. This would be extremely useful if it extends to other models, allowing large models with high pretraining cost to be repurposed with Branching Flows.
One additional observation from this work is that, in the context of Generator Matching, the pre-sampled latent $Z$ is likely underutilized, and Branching Flows provides a template for expanding the space of problems that can be approached in these frameworks.
Further work will explore the performance of Branching Flows in an array of different domains, and more systematically compare the benefits of different components of the process (e.g. branching-only vs deletion-only vs branching & deletion), as well as hyperparameters such as hazard distributions, and the optimal choice of base process for different kinds of problems.
8 Acknowledgments
This project received support from the Swedish Research Council (2024-00390 and 2023-02516) and the Knut and Alice Wallenberg Foundation (2024.0039) to B.M. Key dependencies (e.g. https://github.com/MurrellGroup/Onion.jl) used in this work were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre. AI disclosure: GPT5 was used for initial drafting of some sections, which were rewritten and checked by the authors.
Appendix A Counting and Deletion Process Flow Matching
A.1 Counting Process CTMC
A.1.1 CTMC Equivalence with Counting Event Times
Lemma 1.
Let $T_{1},T_{2},\dots,T_{n}\sim H$ be independent samples from the hazard distribution $H$, with cumulative distribution function $F_{H}(t)$ and density $h(t)$. Define

### v3

@@ Appendix A $Z$ sampling schemes
Here we outline concrete schemes for sampling $Z:=(X_{1}^{+ø},X_{0},\mathcal{T},\mathcal{A})$, starting from $X_{1}$, that we use in our empirical examples below. Some datasets have elements that are naturally grouped (e.g. proteins, grouped by chain), and one can construct $Z$ to respect these groupings, where it makes sense to do so (see section 4.3 for an example). In the case where there is only one group, all elements have the same group index. Further, if any elements are ‘fixed’ (e.g. in the protein example below, by a conditioning mask), then any elements with a fixed element between them in the sequence must belong to a different group.
A.1 Initial Distribution: $X_{0}\sim\mathcal{D}_{0}|(X_{1}=x_{1})$
@@
A.3 Forest: $\mathcal{T}\sim\mathcal{D}_{\mathcal{T}}|(X_{1}^{+ø}=x_{1}^{+ø},X_{0}=x_{0})$
@@
A.4 Internal Anchors: $\mathcal{A^{I}}\sim\mathcal{D}|(X_{1}^{+ø}=x_{1}^{+ø},X_{0}=x_{0},\mathcal{T})$
Anchors control the conditional process along branches, guiding each element so that it would terminate, were $t$ to reach 1, at the anchor. Along terminal branches, the anchors are equal to the elements of $X_{1}^{+ø}$, but along internal branches we have freedom in how we construct them.
Consider a path along one tree from the root to a single leaf. An element drifts, along this path, towards one anchor after another, changing each time a split event occurs. If the anchors change too erratically, the process will be difficult to learn. One principle is to sample anchors that avoid, as much as possible, high-velocity changes.
Since anchors can be sampled conditioned on $X_{1}^{+ø}$, one convenient approach is to specify an anchor merging distribution $p(a_{\text{parent}}|a_{\text{child}_{1}},a_{\text{child}_{2}},\mathcal{T})$ and recursively determine the anchors from leaves to root. This is depicted as $m(\cdot,\cdot)$ in figure 2. Since ordered sequences of continuous values (especially for spatial positions) often exhibit strong autocorrelation with similar values for neighboring elements but discrete values typically do not, the anchor merging schemes can differ depending on whether the state components are continuous or discrete. For our empirical investigations, we use:
  • Continuous ($\mathbb{R}^{n}$): mean of $a_{\text{child}_{1}}$ and $a_{\text{child}_{2}}$, weighted by the number of descendants of each child.
  • Manifold: Geodesic interpolation between $a_{\text{child}_{1}}$ and $a_{\text{child}_{2}}$, weighted by the number of descendants of each child.
@@
Appendix B Counting and Deletion Process Flow Matching
We first describe Counting Flows and Deletion Flows, which we will use to construct Branching Flows.
B.1 Counting Flows
B.1.1 Conditional distributions
We first introduce ‘Counting Flows’: flow matching over a Poisson counting process that generates samples from the space of non-negative integer ‘counts’ via a conditional process that repeatedly increments by 1 to terminate at an observed count $z_{c}$ by $t=1$. This will later be repurposed for Branching Flows to model the counts of ”splitting events” ahead of a node at time $t$, but we present it here in isolation.
Let $H_{\mathrm{inc}}$ be a hazard distribution with no atom at $1$ supported on $[0,1]$, implying that $S_{H_{\mathrm{inc}}}(1)=1-F_{H_{\mathrm{inc}}}(1)=0$. Since $S_{H_{\mathrm{inc}}}(t)$ is the probability of the event not occurring by time $t$, this implies the event occurs almost surely by $t=1$. The hazard rate associated to $H_{\mathrm{inc}}$ is given by
$h_{\mathrm{inc}}(t)=\frac{f_{H_{\mathrm{inc}}}(t)}{1-F_{H_{\mathrm{inc}}}(t)}.$
We also note that since
$\int_{0}^{t}h(s)ds=-\log(1-F_{H_{\mathrm{inc}}}(t)),$
a hazard $h(t)\geq 0$ gives rise to a hazard distribution supported on $[0,1]$ with no atom at $1$ as long as
$\int_{0}^{1}h(t)dt=\infty.$
$Q_{t}^{z_{c}}(x_{t}+1|x_{t})$ $=(z_{c}-x_{t})h_{\mathrm{inc}}(t)I_{\{x_{t}<z_{c}\}},$
and $Q_{t}^{z_{c}}(x_{t}|x_{t})=-Q_{t}^{z_{c}}(x_{t}+1|x_{t}),$ with $Q_{t}^{z_{c}}(\cdot|x_{t})=0$ otherwise.
For any $t$, the increment rate is thus proportional to the number of increments remaining, and the process will almost surely terminate at $z_{c}$.
$D_{R}(R_{t}^{z_{c}},R_{t}^{\theta})=R^{\theta}_{t}-R^{z_{c}}_{t}\log R^{\theta}_{t}$ (2)
is a valid loss under an appropriately scaled time-dependent linear parametrization (billera2025timedependentlossreweighting).
It follows that sampling single-increment independent and identically distributed unordered event times from $H_{\mathrm{inc}}$ corresponds to sampling from the above CTMC, and for a given $t,x_{t},z_{c}$ we can efficiently sample the waiting time until the next increment event.
Lemma B.1.1.
Let $T_{1},T_{2},\dots,T_{n}\sim H$ be independent samples from the hazard distribution $H$, with cumulative distribution function $F_{H}(t)$ and hazard $h(t)$. Define